OpenStudy (anonymous):

Log5(x+2)+log5(2x-1)=2

OpenStudy (anonymous):

\[\log_{5}(x+2) + \log_{5}(2x-1)=2\] As \[\log_{a}b+\log_{a}c=\log_{a}bc\] \[\log_{5}[(x+2)(2x-1)]=2\] \[[(x+2)(2x-1)]=5^{2}\] now simplify.

OpenStudy (anonymous):

Do I solve regularly or distribute the parenthesis to each other